Welland Co. shows the following information on its 2016 income statement: sales $155,500; costs $81,400; ather expenses $4,700; depreclation expense = $10,400; interest expense $7,900; taxes = $17,885; dividends $7,450. In addition, youre told that the fir m issued $3.100 in new equity during 2016 and redeemed $4,900 in outstanding long-term debt a. What is the 2016 operating cash flow? (Do not round Intermedlate calculations and round your answer to the nearest whole number, e.g 32.) Operating cash flow51513 b. What is the 2016 cash flow to creditors? (Do not round intermediate calculations and round your answer to the nearest whole number, .g 32.) Cash flow to creditorsS c. What is the 2016 cash flow to stockholders? (Do not round Intermedlate calculations and round your answer to the nearest whole number .g, 32.) Cash flow to stockholders $ d. If net fixed assets increased by $20,500 during the year, what was the addition to NWC? (Do not round intermediate calculations and round your answer to the nearest whole number, e-g,32.) Addition to net working capital
Answer #1

Cash Flow from Operating activities in 2016 = Sales - ( Costs + Other Expenses + Interest Expenses + Taxes)

= $155,500 - $(81,400 + 4,700 + 7,900 + 17,885)

= $ 43615

Cash Flow to Creditors = Interest Expenses Paid + Redemption of outstanding long term debt

  = $ (7900 + 4900) = $ 12800

Cash Flow to Stockholders = Dividends Paid - New Equity Issued = $ (7450-3100) = $ 4350

(It is assumed that all the costs, other exp, taxes, interest exp and dividends have been actually paid during the year)

________________________________________________

Addition to Net Working Capital (NWC)

Addition to NWC is the incremental (Current Assets - Current Liabilities).

As we are not given any details about both the items, we use an alternative method of calculating

NWC = New Equity and debt issued during the year + Net Profits during the Year - Dividends paid - Debt redeemed during the year - Net Fixed Assets added during the year

=> NWC = 3100 + 33215 - 7450 - 4900 - 10100 = $ 13865

Working:

1. Net Profits during the Year = Cash Flow from operating Activities - Depreciation Exp. = 43615 - 10400 = $ 33215

2. Net Fixed Assets added during the year = New assets purchased - Depreciation for the year

= $ 20500 - $ 10400 = $10100

(The concept is very simple. If you take a balance sheet in your hand, and you hide the current assets and current liabilities, then you can see only equity, debt and Non current assets (fixed assets). So, either you find equity capital plus reserves plus debt minus net fixed assets for the beginning of the year and end of the year, and then find the difference, or you simply calculate the additions made to these line items during the year and find the difference, which has been done above.)

[(Net Profits for the year - dividends paid ) can be called the Additions to Reserves during the year.]